# ChainReaction! Structured Approach with Causal Chains as Intermediate Representations for Improved & Explainable Causal Video Question Answering

### v1

@@ Abstract
Existing Causal-Why Video Question Answering (VideoQA) models often struggle with higher-order reasoning, relying on opaque, monolithic pipelines that entangle video understanding, causal inference, and answer generation. These black-box approaches offer limited interpretability and tend to depend on shallow heuristics. We propose a novel, modular framework that explicitly decouples causal reasoning from answer generation, introducing natural language causal chains as interpretable intermediate representations. Inspired by human cognitive models, these structured cause-effect sequences bridge low-level video content with high-level causal reasoning, enabling transparent and logically coherent inference. Our two-stage architecture comprises a Causal Chain Extractor (CCE) that generates causal chains from video-question pairs, and a Causal Chain-Driven Answerer (CCDA) that produces answers grounded in these chains. To address the lack of annotated reasoning traces, we introduce a scalable method for generating high-quality causal chains from existing datasets using large language models. We also propose CauCo, a new evaluation metric for causality-oriented captioning. Experiments on three large-scale benchmarks demonstrate that our approach not only outperforms state-of-the-art models, but also yields substantial gains in explainability, user trust, and generalization—positioning the CCE as a reusable causal reasoning engine across diverse domains.
@@ 1 Introduction
Understanding the motivations behind human actions is crucial for developing advanced systems capable of nuanced behavior analysis. Human actions are shaped by various factors, including personal experiences, emotions, social contexts, and cultural backgrounds. This complexity demands a comprehensive approach to uncovering the underlying causes of actions. In this context, Causal Video Question Answering (Causal-Why Video QA) requires models not just to recognize events, but to explain why they occurred—demanding higher-order reasoning that surpasses the descriptive QA paradigm (Li, Niu, and Zhang 2022; Xiao et al. 2021; Parmar et al. 2024; Foss et al. 2025).
However, existing Causal-Why Video QA models often draw conclusions from incomplete evidence or rely on shallow heuristics, such as matching action verbs or object nouns in vision-language embedding spaces (Parmar et al. 2024; Rawal et al. 2024; Wei et al. 2023). These models typically entangle video understanding, reasoning, and answer generation into a single monolithic process, which makes their reasoning opaque and error-prone. Moreover, many high-performing VLM models operate as ”black boxes,” offering limited interpretability into their decision-making processes.
Figure 1: Concept. (1) Existing Video ($\mathcal{V}$) Question ($\mathcal{Q}$) Answer ($\mathcal{A}$) approaches through the lens of structural causal models (SCMs), highlighting their monolithic and black-box nature. (2) In contrast, we propose a principled departure from this paradigm: a mediation mechanism based on natural language causal chains ($\mathcal{C}$). We factorize this SCM into two SCMs (3,4)—enabling structured video understanding, reasoning, and inference—leading to superior explainability and performance.
In this work, we make the case that reasoning and answering should be explicitly decoupled and modularized. We introduce a new paradigm in which causal reasoning and answer generation are handled by separate modules that communicate via natural language causal chains (Hanson 1955; Pearl and Mackenzie 2018)—structured sequences of cause-effect events that serve as intermediate reasoning steps (Figure 1). Causal chains provide logically coherent representations that bridge low-level video content with high-level causal understanding (Trabasso and Van Den Broek 1985). To our knowledge, ours is the first approach in Causal-Why VideoQA to explicitly use causal chains as interpretable intermediates, inspired by their role in human cognition and scientific explanation (Gopnik et al. 2004).
@@
Our model consists of two stages: 1) a causal chain extractor (Figure 1(3)), and 2) a causal chain-driven answerer (Figure 1(4)). The causal chain extractor model learns to extract causal chains from video, conditioned on the causal-why questions. The causal chain-driven answerer model learns to generate answers to causal-why questions based on the extracted causal chains. However, a key challenge lies in training causal chain extractor. In particular, we do not have reasoning traces for training causal chain extractor. To tackle this challenge, we develop an approach to generate causal chains from existing VideoQA datasets efficiently.
@@
  • Proposing a structured framework for Causal-Why Video Question Answering that leverages natural language causal chains as intermediate representations to enhance reasoning and transparency.
@@
  • Presenting a novel, scalable and efficient method for generating causal chain annotations using large language models, enabling the augmentation of existing datasets.
@@ 2 Related Work
is a prompting technique that improves large language model performance by decomposing tasks into intermediate reasoning steps. It has proven effective in complex tasks involving arithmetic, logic, and commonsense reasoning, particularly through few-shot prompting with exemplar rationales (Wei et al. 2022). Unlike most CoT-based methods that treat reasoning as latent or emergent, we model reasoning steps—causal chains—as explicit, structured, and learnable variables. This enables supervision, interpretability, and integration with modular architectures, breaking from the monolithic CoT paradigm.
Causal Reasoning and Structural Causal Models (SCMs).
@@
Causality has long been a cornerstone of scientific reasoning, formalized in AI through SCMs and do-calculus (Pearl 2009). Recent work highlights the importance of causal representation learning for improving robustness and generalization in AI systems (Schölkopf et al. 2021). While SCMs and causal inference methods have been applied in contexts like debiasing (Tang et al. 2020) and counterfactual explanations (Narendra et al. 2018), they have rarely been extended to VideoQA. We address this gap by integrating SCMs into a modular framework where causal chains derived from video content support interpretable answer generation. This positions our approach as a novel instantiation of structured causal reasoning in a domain where temporal complexity often hinders transparency.
Causal Chains and Structured Reasoning in AI.
@@
Causal chains are central to human explanations, offering intuitive, step-wise accounts of how and why events unfold (Parmar et al. 2024; Niu et al. 2025). In AI, structured representations like causal graphs have shown promise for improving reasoning and interpretability. Prior efforts, such as (Wang et al. 2024; Ayazoglu et al. 2013), investigate causal interactions in videos or use interventions for action recognition. However, these methods stop short of leveraging causal chains as intermediates for reasoning tasks. Our work formalizes causal chains as explicit bridges between raw video observations and high-level causal understanding, enabling a principled and interpretable approach to Causal-Why Video QA.
Video Causal Reasoning and Temporal Understanding.
A number of recent efforts explore causality in video understanding. For example, (Li et al. 2020; Yang et al. 2021) employ structural models to identify causal interactions in dynamic systems and moment retrieval. However, these do not address question answering. Some methods identify key video segments for question answering (Wei et al. 2023), while others model object interactions over space and time using generative models (Chen et al. 2023). Yet, causal reasoning remains largely implicit or limited to local interactions. Most relevant to our work is (Chen et al. 2025), who propose a dataset and task for identifying event-level causal diagrams in videos. Their model infers why a final event occurred but does not generate full causal chains or use them for causal question answering. By contrast, we extract event-level causal chains to directly power a dedicated inference module. Zang et al. (Zang et al. 2023) explore video-text causal links, while Su et al. (Su et al. 2023) generate causal questions from captions, but they do not model causal chains or explicitly decouple reasoning from answering.
Vision-Language Models and Their Limitations in Causality.
@@
Recent Vision-Language Models (VLMs) such as VideoLLaMA (Zhang, Li, and Bing 2023), VideoChat2 (Li et al. 2024), and VILA (Lin et al. 2024) have advanced video QA and captioning by leveraging large-scale video-text pretraining. However, these models focus on recognizing and describing events, not explaining them. Their monolithic architectures lack interpretability and often rely on shallow correlations. Our approach complements these models by introducing causal chains as structured reasoning intermediates—enhancing both performance and transparency in Causal-Why Video QA.
Figure 2: Causal chain construction for SFT. (1) Human annotators of base datasets intuitively and implicitly make use of causal chains when writing correct answers. (2) We propose to recover these causal chains with the help of LLM using questions and correct gold answers. (3,4) Our robust causal chain generation and verification pipeline.
@@ 3 Causal Chain Construction for SFT
In this section, we propose a novel, scalable and efficient methodology for the construction of causal chains to be used for supervised finetuning (SFT) of our causal chain extractor (presented in later sections). Our goal is to generate causal chains that accurately reflect the reasoning behind human-written answers to video-based questions. Since no existing VideoQA dataset includes causal chain annotations to support our approach, we collect causal chains for three challenging causal video QA datasets: 1) NextQA (Xiao et al. 2021), 2) CausalVidQA (Li, Niu, and Zhang 2022), and 3) CausalChaos! (Parmar et al. 2024). We refer to these three datasets as the base (or source) datasets. We focus on Causal-Why QA in these base datasets. In base datasets, each sample is a paired triplet of {Video ($\mathcal{V}$), Question ($\mathcal{Q}$), Answer ($\mathcal{A}$)}. We have provided summaries of these three base datasets in the Appendix.
Preliminary I.
@@
Human annotators of the base dataset watch the video ($\mathcal{V}$), then formulate a question ($\mathcal{Q}$) related to video; and then finally, meticulously write the correct, gold answer ($\mathcal{A}$) to the question. Further, the annotations are rigorously cross-annotator verified, resulting in multihuman agreement. Human annotators intuitively and implicitly make use of causal chains ($\mathcal{C}$) in the process of writing answers. This process is illustrated in Figure 2(1). As a result, the detailed QA pairs of our base datasets contain causal chains embedded within them.
Preliminary II.
@@
Given the questions and the corresponding correct, gold answers, LLMs have been demonstrated to accurately recover the intermediate reasoning steps or causal chains. Illustrated in Figure 2(2).
Corollary I.
@@
Based on preliminaries I and II, we propose that if provided with the questions and the corresponding human written correct, gold answers from the base datasets, then powerful oracle LLMs (e.g., GPT4o, DeepSeek, Gemini) can reliably recover the reasoning steps or causal chains.
Causal Chain Construction Methodology.
@@
Following Corollary I, we propose a robust methodology for causal chain generation and verification in the following.
  1. 1.
@@
    Causal Chain Generation Process: We leverage a powerful LLM (GPT4o). We refer to it as Oracle LLM. We prompt the Oracle LLM with question ($\mathcal{Q}$) and the corresponding human written correct, gold answer ($\mathcal{A}$) and ask it to generate a causal chain ($\mathcal{C}$) in natural language in a specified format. Specifically, we instruct the LLM to return causal chains in a structured format: [Event A] → [Event B] → [Event C]... [^1] Note that we do not train the LLMs to generate causal chains; rather, directly capitalize on their inherent knowledge, commonsense reasoning & synthesis capabilities to derive them. Full prompt provided in the Appendix. Illustrated in Figure 2(3).
  2. 2.
@@
    1. (a)
      Programmatic Validation: Programs are developed for checking proper structure, formatting, completeness and length (constraint: $\leq$ 10 events) of the generated causal chains.
    2. (b)
@@
    3. (c)
      Manual Verification against $\mathcal{V}$: Human verifiers are provided with full access to the corresponding video $\mathcal{V}$, the question $\mathcal{Q}$, and the correct gold answer $\mathcal{A}$ to ensure proper grounding of the causal chain in the video content. Chains passing the first two stages are verified by humans for logical coherence, relevance, and consistency with $\mathcal{Q}$ and $\mathcal{A}$ pair and grounded in video ($\mathcal{V}$). Human verifiers are computer science graduates and well-versed with the task and literature. Human verifiers are allowed to manually add missing details to the causal chains. In addition, human verifiers are trained to identify hallucinations (e.g., events not observable or implied in the video) during manual verification. Chains containing any such hallucinations are rejected and regenerated.
    If the chain fails at any stage, it is regenerated, until it passes all the checks. Finally, 1000 samples belonging to each base dataset are manually reviewed by the authors. On average, over 95% of the chains passed the author verification, ensuring high-quality annotations.
Our methodology allowed to generate consistent and reliable causal chains with respect to the video context. In total, we collected causal chains for 46024 samples across three base datasets. Further statistics provided in Appendix.
@@ 4 Approach
Theoretical Foundation and Motivation
Our work is grounded in the principles of Structural Causal Models (SCMs) (Pearl 2009) and Chain-of-Thought (CoT) (Wei et al. 2022) reasoning. We decompose the VideoQA task into two explicit stages: 1) causal chain extraction; and 2) causal chain-driven answering. Please refer to Figure 1. From an SCM perspective, we introduce causal chains as structured intermediate representations. While many existing VideoQA approaches can be abstracted as an undifferentiated model where video ($\mathcal{V}$) and question ($\mathcal{Q}$) jointly influence the answer ($\mathcal{A}$) as ($\mathcal{V}\rightarrow\mathcal{A}\leftarrow\mathcal{Q}$), we propose a structured mediation via a causal chain variable ($\mathcal{C}$), yielding the following SCM: $\mathcal{V}\rightarrow\mathcal{C}\rightarrow\mathcal{A},\mathcal{Q}\rightarrow\mathcal{C}\rightarrow\mathcal{A}$. Then, we factorize this SCM into: $\mathcal{V}\rightarrow\mathcal{C}\leftarrow\mathcal{Q}$ (this becomes our Causal Chain Extractor—Figure 1(3)) and $\mathcal{Q}\rightarrow\mathcal{A}\leftarrow\mathcal{C}$ (this becomes our Causal Chain-driven Answerer—Figure 1(4)). Causal Chain Extractor module would first extract $\mathcal{C}$, and pass it to the Causal Chain-Driven Answerer module. This structured factorization aligns with CoT principles by making reasoning steps explicit, and reflects SCM modularity by separating causal understanding from answer generation. As a result, our approach enables:
  • Focused processing: Unlike typical VideoQA approaches, the proposed structured approach allows more focused processing at each stage, with the outputs of each stage explicitly made available to the subsequent one. By handling smaller, more focused chunks, there is a lesser risk of overlooking important reasoning steps or drawing incorrect conclusions unlike existing approaches. This approach draws inspiration from the Chain-of-Thought philosophy, which we adapt into a structured model.
  • Improved video understanding: Modeling cause-and-effect relationships explicitly in causal chains is a dense prediction task. With detailed supervision, vision models learn to better able to capture rich video content. In contrast, typical VideoQA models are trained with a single label (the correct answer), which may be insufficient to fully characterize the video’s complexity (Bi et al. 2025).

### v2

# ChainReaction: Causal Chain-Guided Reasoning for Modular and Explainable Causal-Why Video Question Answering

@@ Abstract
Existing Causal-Why Video Question Answering (VideoQA) models often struggle with higher-order reasoning, relying on opaque, monolithic pipelines that entangle video understanding, causal inference, and answer generation. These black-box approaches offer limited interpretability and tend to depend on shallow heuristics. We propose a novel, modular paradigm that explicitly decouples causal reasoning from answer generation, introducing natural language causal chains as interpretable intermediate representations. Inspired by human cognitive models, these structured cause-effect sequences bridge low-level video content with high-level causal reasoning, enabling transparent and logically coherent inference. Our two-stage architecture comprises a Causal Chain Extractor (CCE) that generates causal chains from video-question pairs, and a Causal Chain-Driven Answerer (CCDA) that derives answers grounded in these chains. To address the lack of annotated reasoning traces, we introduce a scalable method for generating accurate causal chains from existing datasets. We construct human verified causal chains for 46K samples. We also propose CauCo, a new evaluation metric for causality-oriented captioning. Experiments on three large-scale benchmarks demonstrate that our approach not only outperforms state-of-the-art models, but also yields substantial gains in explainability, user trust, and generalization—positioning the CCE as a reusable causal reasoning engine across diverse domains.
@@ 1 Introduction
Understanding the motivations behind human actions is crucial for developing advanced systems for nuanced behavior analysis. Human actions are shaped by factors such as personal experience, emotion, social context, and culture. This complexity requires uncovering underlying causes. In this context, Causal Video Question Answering (Causal-Why VideoQA) asks models not only to recognize events but to explain why they occur—demanding higher-order reasoning beyond descriptive QA [causalvidqa, nextqa, causalchaos, foss2025causalvqa].
Existing Causal-Why VideoQA models often reason from incomplete evidence or rely on shallow heuristics (e.g., matching action verbs or object nouns in vision-language embedding spaces [causalchaos, rawal2024dissecting, wei2023visual]). These models entangle video understanding, reasoning, and answer generation into one monolithic process, making their reasoning opaque and error-prone. Many high-performing vision language models (VLMs) also operate as black boxes, offering limited interpretability into their decisions.
Figure 1: (Top) Concept. (1) Existing Video ($\mathcal{V}$) Question ($\mathcal{Q}$) Answer ($\mathcal{A}$) approaches through the lens of structural causal models (SCMs), highlighting their monolithic and black-box nature. (2) In contrast, we propose a principled departure from this paradigm: leveraging the Causal Reasoning Trace ($\mathcal{C}$), a structured intermediate representation based on natural language causal chains. We factorize this SCM into two SCMs (3,4)—enabling structured video understanding, reasoning, and inference—leading to superior explainability and performance. (Bottom) Example. Please zoom in for the best view.
In this work, we make the case that reasoning and answering should be explicitly decoupled and modularized. We introduce a new paradigm in which causal reasoning and answer generation are handled by separate modules that communicate via natural language causal chains [hanson1955causal, pearl2018book]—structured sequences of cause-effect events that serve as intermediate reasoning steps (Figure 1). Causal chains or causal reasoning traces provide a logically coherent bridge between low-level video content and high-level causal understanding [trabasso1985causal]. Formulated as natural-language sequences, they capture the linear, observable steps linking a cause (e.g., a character’s intention) to its effect, ensuring that answers are grounded in the video’s causal progression. To our knowledge, ours is the first approach in Causal-Why VideoQA to explicitly use causal chains as interpretable intermediates, inspired by their role in human cognition and scientific explanation [gopnik2004theory].
@@
Our model consists of two stages: 1) a causal chain extractor—CCE—(Figure 1(3)), and 2) a causal chain-driven answerer—CCDA—(Figure 1(4)). The CCE model learns to extract causal chains from video, conditioned on the causal-why questions. The CCDA model learns to generate answers to causal-why questions based on the extracted causal chains. Distinct from general causal inference models that reason over partially observed or hypothetical systems, the CCE operates in fully observed video settings, where the causal process has already unfolded. After observing the complete video, the CCE identifies the actualized, linear sequence of events that led to the outcome—making the task descriptive and post-hoc rather than inferential, and emphasizing interpretability and causal fidelity within the VideoQA framework. However, a key challenge lies in training CCE. In particular, there are no datasets containing reasoning traces for training CCE. To tackle this challenge, we develop an approach to generate causal chains from existing VideoQA datasets efficiently.
@@
  • Proposing a structured paradigm for Causal-Why Video Question Answering that leverages natural language causal chains as intermediate representations to enhance reasoning and transparency.
@@
  • Introducing a human-in-the-loop framework that uses large language models to propose causal chain drafts, which are subsequently verified and finalized by human annotators, yielding high-quality causal reasoning data.
@@ 2 Related Work
Chain-of-Thought (CoT) is a prompting technique that improves LLMs by decomposing tasks into intermediate reasoning steps. It is effective in arithmetic, logic, and commonsense tasks, especially via few-shot prompting with exemplars [wei2022chain]. Unlike CoT methods that treat reasoning as emergent, we model reasoning steps—causal chains—as explicit, structured variables. This enables supervision, interpretability, and integration with modular architectures, breaking from the monolithic CoT paradigm.
@@
Causal Reasoning and Structural Causal Models (SCMs). Causality is foundational in scientific reasoning and is formalized in AI through SCMs and do-calculus [pearl2009causality]. Recent work emphasizes causal representation learning for robustness and generalization [scholkopf2021toward], and SCM-based methods have been used for tasks like debiasing [tang2020unbiased] and counterfactual explanations [narendra2018explaining]. However, such approaches are rarely applied to VideoQA. We address this gap by integrating SCMs into a modular framework where video-derived causal chains support interpretable answer generation—bringing structured causal reasoning to a domain where temporal complexity often obscures transparency.
@@
Causal Chains and Structured Reasoning in AI. Causal chains provide intuitive, stepwise explanations of how events unfold [causalchaos, niu2025r]. In AI, structured representations like causal graphs improve reasoning and interpretability, and prior work has explored causal interactions in videos or intervention-based action recognition [wang2024modeling, ayazoglu2013finding]. However, these approaches do not use causal chains as intermediates for reasoning. We instead formalize causal chains as explicit bridges between video observations and high-level causal understanding, enabling a principled and interpretable framework for Causal-Why Video QA.
Video Causal Reasoning & Temporal Understanding. Recent work explores causality in video understanding—e.g., structural models for causal interactions and moment retrieval [li2020causal, yang2021deconfounded]—but not for QA. Other methods select key segments for QA [wei2023visual] or model object interactions generatively [chen2023llcp], yet their causal reasoning remains implicit or localized. Chen et al. [chen2025mecd] introduce event-level causal diagram annotation, but their models explain a final event rather than generate causal chains or use them for QA. In contrast, we extract event-level causal chains and leverage them in a dedicated inference module. Zang et al. [zang2023discovering] study video–text causal links, and Su et al. [su2023language] generate causal questions from captions, but neither models causal chains or decouples reasoning from answering.
@@
Vision-Language Models (VLMs) & their Limitations in Causality. Recent VLMs (e.g., VideoLLaMA [videollama], VideoChat2 [videochat2], VILA [vila], etc.) advance VideoQA & captioning through largescale video–text pretraining, but primarily recognize and describe events rather than explain them. Their monolithic architectures offer limited interpretability & often rely on shallow correlations. Our method complements them by introducing causal chains as structured reasoning intermediates, improving both performance & transparency in Causal-Why Video QA.
@@ 3 Causal Chain Construction for SFT
We propose a novel, scalable and efficient methodology to construct causal chains to be used for supervised finetuning (SFT) of our causal chain extractor (Section 4). We generate causal chains that accurately reflect the reasoning behind human-written answers to video-based questions.
@@
Base datasets: since no existing VideoQA dataset includes causal chain annotations to support our approach, we collect causal chains for three challenging causal video QA datasets: 1) NextQA [nextqa], 2) CausalVidQA [causalvidqa], and 3) CausalChaos! [causalchaos]. We refer to these three datasets as the base (or source) datasets. We focus on Causal-Why QA in these base datasets. In base datasets, each sample is a paired triplet of {Video ($\mathcal{V}$), Question ($\mathcal{Q}$), Answer ($\mathcal{A}$)}.
Causality defined in the base datasets: the causal-why questions and correct answers in base datasets are human-authored, as a result the notion of cause or causality is well defined in the data itself. All the annotations are rigorously cross-annotator verified, resulting in multihuman agreement on causality. Our goal is to learn to reason through these causal relations. We have provided further details on these three base datasets in the Appendix.
@@
Preliminary I. Human annotators watch the video ($\mathcal{V}$), formulate a question ($\mathcal{Q}$) about it, and write the correct/gold answer ($\mathcal{A}$). The annotations are cross-verified, ensuring multi-annotator agreement. Annotators implicitly use causal chains ($\mathcal{C}$) when writing answers, as illustrated in Figure 2(1). Thus, the detailed QA pairs of our base datasets contain causal chains embedded within them.
@@
Preliminary II. Given the questions and the corresponding correct, gold answers, LLMs have been demonstrated to accurately recover the intermediate reasoning steps or causal chains. Illustrated in Figure 2(2).
@@
Corollary I. Based on preliminaries I and II, we propose that if provided with the questions and the corresponding human written correct, gold answers from the base datasets, then powerful oracle LLMs (e.g., [gpt4o, deepseekvl2, gemini]) can reliably recover the reasoning steps or causal chains.
Figure 2: Causal chain construction for SFT. (1) Human annotators of base datasets intuitively and implicitly make use of causal chains when writing correct answers. (2) We propose to recover these causal chains with the help of LLM using questions and correct gold answers. (3,4) Our robust causal chain generation and manual verification and video grounding check pipeline.
@@
Causal Chain Construction Methodology. Following Corollary I, we propose a robust methodology for causal chain generation and verification in the following.
  1. wide
@@
    Causal Chain Draft Generation Process: We leverage a powerful LLM (GPT4o). We refer to it as Oracle LLM. We prompt the Oracle LLM with question ($\mathcal{Q}$) and the corresponding human written correct, gold answer ($\mathcal{A}$) and ask it to generate a causal chain ($\mathcal{C}$) in natural language in a specified format. Specifically, we instruct the LLM to return causal chains in a structured format: [Event A] → [Event B] → [Event C]... [^1] Note that we do not train the LLMs to generate causal chains; rather, directly capitalize on their inherent knowledge, commonsense reasoning & synthesis capabilities to derive them. Full prompt provided in the Appendix. Illustrated in Figure 2(3). Oracle LLM is not used as end-task solvers but as structured reasoning components within a rigorously controlled pipeline. Their role is analogous to that of annotator assistants—providing drafts that undergo multi-layer human validation to ensure reliability and reproducibility.
  2. wiide
@@
    1. wide
      Programmatic Validation: Programs check generated causal chains for structure, format, completeness, and length ($\leq$10 events).
    2. wiide
@@
    3. wiiide
      Manual Correction & Verification against $\mathcal{V}$: Human verifiers receive the video $\mathcal{V}$, question $\mathcal{Q}$, and correct gold answer $\mathcal{A}$ to ensure proper grounding of the causal chain in the video content. Chains passing the first two stages are checked for logical coherence, relevance, and consistency with $\mathcal{Q}$, $\mathcal{A}$, and $\mathcal{V}$. Verifiers—computer science graduates familiar with the task—may add missing details to causal chains and flag hallucinations (e.g., events not observable or implied in the video). Chains with hallucinations or other failures are regenerated until they pass all checks.
Only the chains that have passed all the checks are considered for the final dataset. Finally, 1000 samples belonging to each base dataset are manually reviewed by the authors. Over 95% of the chains passed the author verification, ensuring accurate causal chain annotations. Our method produced reliable, accurate causal chains grounded in the video context. In total, we constructed human-verified causal chains for 46,024 samples across three datasets. Further stats provided in the Appendix. Dataset will be released for reproducibility and future research.
@@ 4 Approach
4.1 Overview and Motivation
Conceptually, our work is grounded in the principles of Structural Causal Models (SCMs) [pearl2009causality] and Chain-of-Thought (CoT) [wei2022chain] reasoning. We decompose the VideoQA task into two explicit stages: 1) causal chain extraction; and 2) causal chain-driven answering. Please refer to Figure 1. From an SCM perspective, we introduce causal chains as structured intermediate representations. While many existing VideoQA approaches can be abstracted as an undifferentiated model where video ($\mathcal{V}$) and question ($\mathcal{Q}$) jointly influence the answer ($\mathcal{A}$) as ($\mathcal{V}\rightarrow\mathcal{A}\leftarrow\mathcal{Q}$), we propose a structured reasoning via a Causal Reasoning Trace ($\mathcal{C}$), yielding the following SCM: $\mathcal{V}\rightarrow\mathcal{C}\rightarrow\mathcal{A},\mathcal{Q}\rightarrow\mathcal{C}\rightarrow\mathcal{A}$. Then, we factorize this SCM into: $\mathcal{V}\rightarrow\mathcal{C}\leftarrow\mathcal{Q}$ (this becomes our Causal Chain ExtractorFigure 1(3)) and $\mathcal{Q}\rightarrow\mathcal{A}\leftarrow\mathcal{C}$ (this becomes our Causal Chain-driven AnswererFigure 1(4)). Causal Chain Extractor module would first extract $\mathcal{C}$, and pass it to the Causal Chain-Driven Answerer module.
We assume access to fully observed videos, where the causal process leading to an outcome manifests as a temporally ordered sequence of events—a single realized traversal through the underlying causal graph. Once this process is complete, the relevant causal information resides entirely within this factual path. Thus, for Causal-Why VideoQA, the causal chain can be modeled as a linear sequence of observed cause–effect events, providing a complete & interpretable representation of the reasoning process. Detailed treatment of the proposed use of linear causal chains is provided in the Appendix. The structured factorization aligns with CoT principles by making reasoning steps explicit, & reflects SCM modularity by separating causal understanding from answering. As a result, our approach enables:
  • Focused processing: Unlike existing VideoQA methods, our structured approach enables focused processing, with each stage’s output passed to the next. Processing smaller chunks reduces the risk of missing reasoning steps or making incorrect inferences. It is inspired by the Chain-of-Thought philosophy, adapted into a structured model.
  • Improved video understanding: Modeling cause-and-effect relationships explicitly in causal chains is a dense prediction task. With detailed supervision, vision models learn to better capture rich video content. In contrast, typical VideoQA models are trained with a single label, which may be insufficient to capture the video’s complexity [bi2025reasoning].